OpenStudy (anonymous):

sin(a)-cot(a)+cot(2a)=

OpenStudy (anonymous):

sina-cosec2a

OpenStudy (anonymous):

sina-cota+cot2a=sina-(cosa/sina)+(cos2a/sin2a) =(2sin^2acosa-2cos^a+1-2sin^2a)/sin2a =(2sin^2acosa-2(cos^2a+sin^2a)+1)/sin2a =(2sin^2acosa-1)/sin2a =sina-(1/sin2a) =sina-cosec2a